Given GDP at market price=12000, depreciation = 1500,subsidies =300
Calculate NDP at factor cost
1
Expert's answer
2019-04-10T09:30:27-0400

Solution:


"NDPfc=NDPmp - D"

where, NDPfc - Net Domestic product at factor cost, NDPmp - Net Domestic product at market prices, D - depreciations,


"NDPmp = GDPmp + S"

where, GDPmp - Gross Domestic product at market prices, S - subsidies.

In this way:


"NDPmp = 12000 + 300 = 12300"

"NDPfc = 12300 - 1500 = 10800"

Answer: NDP at factor cost is 10800